OpenStudy (anonymous):

How do I create a quadratic equation for x=9 and x=-1 People keep trying to explain it to me but I don't remember half of what they're talking about?

OpenStudy (anonymous):

Any quadratic formula (x-a)(x-b). Is verified, that the roots are x = a and x = b. So if you want a specific equation just plug em in!

OpenStudy (anonymous):

lets look at it this way. suppose you wanted to solve \[x^2+8x-9=0\] first you try to factor. and you can factor by as \[(x-9)(x+1)=0\] so then you would say \[x-9=0,x=9\] or \[x+1=0,x=-1\]

OpenStudy (anonymous):

so you know the solutions are 9 and -1, because first you factored and then you found them

OpenStudy (anonymous):

now we are just working backwards. you know the solutions are 9 and -1. then must have come from \[x=9, x=-1\] so \[x-9=0, x+1=0\] and the "factored form" must have been \[(x-9)(x+1)=0\]

OpenStudy (anonymous):

if we multiply this out we get \[x^2+8x-1=0\] and that is the answer you are looking for

OpenStudy (anonymous):

(x-9)(x+1)=0 now x-9=0 and x=0 or x+1=0 and x= -1 x^2+8x-1=0 So if I write it out like this, does it make sense?

OpenStudy (anonymous):

or am I skipping important steps?

OpenStudy (anonymous):

all you have to state is that, for x=9 and x=-1 to be solutions, the factored expression will be (x-9)(x+1)=0. then, multiplying it out, you get x^2-8x-9=0

OpenStudy (anonymous):

satellite73's equation is incorrectly multiplied

OpenStudy (anonymous):

How would it look if I multiplied it out?

OpenStudy (anonymous):

(x-9)*(x+1) just FOIL and youll get x^2-9x+x-9=0 combine like terms and you get x^2-8x-9=0

OpenStudy (anonymous):

I know this has go to be a stupid question, but what's FOIL again?

OpenStudy (anonymous):

I just has major issues with my short term memory. It's super frustrating.

OpenStudy (anonymous):

it stands for "First, Outer, Inner, Last", it is just how expanding a polynomial is described

OpenStudy (anonymous):

have*